OpenStudy (anonymous):

-5x+y=0 x+y=6 So that gives me 5x+y=6

OpenStudy (anonymous):

5x=6 6/5= y=1.2

OpenStudy (nurali):

-5x+y=0---(1) x+y=6----(2) subtract (1) and (2) -6x=-6 x=1 put in equation (2) 1+y=6 y=6-1 y=5

OpenStudy (anonymous):

thanks Nurali!

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

I had it all wrong :P